OpenStudy (anonymous):

y = cos x; [0, 4π]; a = 1/2

OpenStudy (ghazi):

what are you supposed to do here?

OpenStudy (anonymous):

the question says... find all the values of x such that for the real number a

OpenStudy (anonymous):

and you have to do it in respect to [0,4pi]

OpenStudy (ghazi):

now your question says find x at which y=1/2 that means cosx=1/2 x=60, and cos is positive in fourth quadrant 240 degrees , pi/3 and 4pi/3

OpenStudy (ghazi):

now pi/3 and 4pi/3 is in the range of (0,2pi) now next two will be 5pi/3 and 3pi

OpenStudy (anonymous):

i can't get it right i typed those in and it keeps saying wrong

OpenStudy (ghazi):

solution is \[\frac{ \pi }{ 3 },\frac{ 4 \pi }{ 3 },\frac{ 5 \pi }{ 3 },3 \pi\]

OpenStudy (anonymous):

it says… y=a and then what does x equal?

OpenStudy (anonymous):

its not working

OpenStudy (ghazi):

if y=a then y=1/2 because a=1/2 therefore cos x= 1/2 and x= pi/3

OpenStudy (ghazi):

cos 60 = 1/2

OpenStudy (anonymous):

i know thats how i did it but its wrong

OpenStudy (ghazi):

then i guess solution is wrong, do you have options?

OpenStudy (anonymous):

no its open ended

OpenStudy (ghazi):

then dude x= pi/3 is the answer and all those that i mentioned above in the interval (0, 4pi)

OpenStudy (anonymous):

tried it

OpenStudy (ghazi):

then there is problem with solution i mean there is some error

OpenStudy (anonymous):

i couldn't get any of them for the problems that were like this

OpenStudy (ghazi):

try in radians, degrees and grads may be one of them will fit in